OpenStudy (anonymous):

if f(x) =5x-7 and g(x)= square root (x-2), which statement is true -5 is in the domain of ƒ º g. -5 is not in the domain of ƒ º g.

OpenStudy (anonymous):

help please.. i dont get it

OpenStudy (aum):

ƒ º g = f(g(x)) Before even actually finding what f(g(x)) is we can say the domain of the composite function should consider the domain of the inner function g(x). What is the domain of g(x) ?

OpenStudy (anonymous):

given f(x) =square root of 8x and g(x) = 1/x-4. which value is in the domain of f g?

OpenStudy (anonymous):

-9 0 4 6

OpenStudy (anonymous):

will you help me please

OpenStudy (aum):

Can you answer my earlier question? \(g(x) = \sqrt{x-1}\) What is the domain of g(x) ?

OpenStudy (anonymous):

i already got that answer